Using dimensional analysis, verify the correctness of
1. v = 2gh and
2. v = ut/2 + at^2
1
Expert's answer
2019-06-10T09:50:13-0400

In dimensional analysis, an equation is correct if the dimension of its left-hand side (LHS) and right-hand side (RHS) are equal, that is


"[LHS] = [RHS]."

The dimension of variables:

final velocity


"v = [LT^{-1}],"

initial velocity

"u=[LT^{-1}],"

acceleration

"g=[LT^{-2}],"

height

"h=[L],"


time

"t=[T],"

number 2 has no dimension, where L is length and T is time.


Solving the problems.


1. For a given expression 

"v = 2gh"


true that


"LHS = v, [LHS] = [LT^{-1}];""RHS = 2gh, [RHS] = [LT^{-2}][L] = [L^2 T^{-2}];""[LT^{-1}]\neq[L^2 T^{-2}],"


that is

"[LHS]\neq[RHS]."

Last inequality means that expression isn`t correct.


2. 


"v = \frac{ut}{2} + at^2;""LHS = v, [LHS] = [LT^{-1}];""RHS = \frac{ut}{2} + at^2,""[RHS] = [L T^{-1}][T]+[L T^{-2}][T^2] = [L]+[L] = [L];""[LHS]\neq[RHS],"

therefore


"[LT^{-1}]\neq[L],"


so, the equation is false.